OpenStudy (anonymous):

Find the critical points for f(t) = t/(1 + t^2)

OpenStudy (anonymous):

Can you do the derivative of the equation? Set that equal to zero.

OpenStudy (anonymous):

Also identify points that will make the derivative undefined.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Get your complete answer?

OpenStudy (anonymous):

I am stuck on finding the crtitical points from (1-2t^2(1+t^2)^-1). I know that the factored out piece (1 + t^2)^-1 is where the derivative is undefined...

OpenStudy (anonymous):

Let me check something. That isn't the derivative that I ended up with.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

You should end up with (1-t^2)/[(1+t^2)^2\

OpenStudy (anonymous):

Then your critical points should be pretty easy to find from there.

OpenStudy (anonymous):

Was your f'(t) = (1+t^2)^-1 - 2t^2(1+t^2)^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I am not sure how you came to your derivative.

OpenStudy (anonymous):

\[f(t) = t/(1+t ^{2})\] define \[g(t) = t\] \[h(t) = 1 + t ^{2}\] Quotient Rule: \[f'(x) = [h(x)g'(x) - g(x)h'(x)]/[h(x)]^{2}\]

OpenStudy (anonymous):

You can also use the product rule, of course. This is just an easy way to annotate online.

OpenStudy (anonymous):

If you use the product rule, remember the chain rule with the negative exponent.

OpenStudy (anonymous):

I got \[f'(x) = ((1+t^2)-2t^2)/(1+t^2)^2\]

OpenStudy (anonymous):

Correct. Simplify the numerator.

OpenStudy (anonymous):

So now I have \[-2t^2/(1+t^2)\]

OpenStudy (anonymous):

Not quite

OpenStudy (anonymous):

Oh! I know what I did-- the actual answer is \[(1-t^2)/(1+t^2)^2\]

OpenStudy (anonymous):

There ya go

OpenStudy (anonymous):

So, what values of t make this 0 or undefined?

OpenStudy (anonymous):

No value can make this undefined because t is squared in the denominator...

OpenStudy (anonymous):

No real value, right. i will, if you are allowed complex solutions. I'm guessing not, however.

OpenStudy (anonymous):

I can't multiply the denominator on each side to get rid of it, right?

OpenStudy (anonymous):

I can graph it...

OpenStudy (anonymous):

You can, but it isn't needed. You have done all the steps you need.

OpenStudy (anonymous):

But the critical points are when I set f' equal to zero...

OpenStudy (anonymous):

A number line would be unrealistic...

OpenStudy (anonymous):

The derivative is 0 when \[1 - t ^{2} = 0\]

OpenStudy (anonymous):

Oh---we only set the numerator equal to zero? Does that apply for all fraction derivatives...because I know that if I set the denominator equal to zero, it would show the values where the function is undefined.

OpenStudy (anonymous):

Right. If the numerator is 0, the fraction is 0. If the denominator is 0, the fraction is undefined. There are some interesting things that happen if both are 0, but you'll learn those later :D

OpenStudy (anonymous):

So I have my critical points as +/- 1. This means that one of them can be a global max/min when I test the critical points.

OpenStudy (anonymous):

They can be global, local, or neither.

OpenStudy (anonymous):

Thanks so much for helping me and staying online this entire time! I really appreciate it!

OpenStudy (anonymous):

My pleasure